OpenStudy (anonymous):

6 review questions help please? medal!

OpenStudy (anonymous):

1. What is the value of h? a. 4 b. 8 sqrt 3 c. 16 d. 8 sqrt 2

OpenStudy (whpalmer4):

Okay, you have an isosceles right triangle here, because both of the angles are the same. What does that imply about the associated sides?

OpenStudy (anonymous):

They have 2 equal sides and 2 equal angles?

OpenStudy (whpalmer4):

Right. So what is the length of the other leg of the triangle?

OpenStudy (anonymous):

4 ?

OpenStudy (whpalmer4):

 What is the value of ???

OpenStudy (anonymous):

I'm still not sure .. /:

OpenStudy (whpalmer4):

come on, we said this was an isosceles triangle, which means two equal angles and two equal sides!

OpenStudy (whpalmer4):

Doesn't that mean ??? has to be 8?

OpenStudy (anonymous):

why would it be 8 ?

OpenStudy (anonymous):

sorry I feel stupid >_<

OpenStudy (whpalmer4):

because it is an isosceles triangle. One of the sides is 8. Another one must be 8 as well, and it can't be the hypotenuse, because the hypotenuse is the longest side.

OpenStudy (whpalmer4):

it's an isosceles triangle because it has two equal angles. You see that part, right?

OpenStudy (anonymous):

ohhhhh ok I understand

OpenStudy (whpalmer4):

So we can update our drawing:  If the two sides labeled 8 are \(a=8,b=8\) respectively, what is the length of the hypotenuse, \(c\)? Use the Pythagorean theorem

OpenStudy (anonymous):

k hold on

OpenStudy (anonymous):

its a^2+b^2=c^2 do I add the 45 degrees together to get c? or is the 8?

OpenStudy (whpalmer4):

The Pythagorean theorem relates only to the lengths of the sides. Angles don't matter (except, of course, that one of them must be 90 degrees for it to be a right triangle).

OpenStudy (whpalmer4):

\[a^2+b^2=c^2\]\[8^2+8^2=c^2\]\[2*8^2 = c^2\]Take the square root of both sides:\[\sqrt{2*8^2} = \sqrt{c^2}\]Simplify\[8\sqrt{2} = c\] If you have a \(45^\circ/45^\circ/90^\circ\) triangle such as this one, the side lengths are always in the ratio \(1:1:\sqrt{2}\)

OpenStudy (anonymous):

is that the way you would do it for this kind of triangle too ?

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

I don't understand your question.

OpenStudy (anonymous):

Use the Pythagorean theorem

OpenStudy (whpalmer4):

Well, yes, if you know 2 sides in a right triangle, you can use the PT to find the 3rd side.

OpenStudy (anonymous):

okay ty c:

OpenStudy (whpalmer4):

you're welcome